OpenStudy (anonymous):

simplify completely x^2 -15x + 36/5x - 60

OpenStudy (anonymous):

factor each of the expressions.

OpenStudy (anonymous):

36/5-15

OpenStudy (anonymous):

= 36/5 -75/5 = -39/5

OpenStudy (anonymous):

 Thx!